# A systematic approach to $\ell$-loop planar integrands from the classical equation of motion

### v3

@@ Abstract
In this letter, we present a recursive method for $\ell$-loop planar integrands in colored quantum field theories. We start with the classical equation of motion and then pick out the comb component, which will help us to define the loop kernels. Then we construct the $\ell$-loop integrands based on some recursion rules for the $\ell$-loop kernels. Finally, we reach a recursion formula for the $\ell$-loop planar integrands. Our method can be easily generalized to general quantum field theories, even non-Lagrangian theories, to obtain the planar part of the whole $\ell$-loop integrands.
@@ I Introduction
Most of the development on the analytic side comes from the success of the on-shell methods, see [1, 2] for some reviews. Besides the on-shell method, off-shell methods have also been developed in various fields [3]. Off-shell objects have a richer structure and can help us know more about different theories. In addition, off-shell objects are very useful for generating loop integrands or computing amplitudes recursively. One of the most successful off-shell objects is the Berends-Giele (BG) current [4]. The BG current, which has only one off-shell leg, has many applications in different cases [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Hence, to develop off-shell techniques is very important both in theory and practical calculations.
@@
In this letter, we want to develop techniques based on the multi-particle solutions. We start with the equation of motion of the bi-adjoint theory, and finally reach the whole $\ell$-loop planar integrand results  [^1] Planar integrands correspond to the partial amplitudes associated with a single-trace structure in the color-trace decomposition [1].. This systematic approach to $\ell$-loop planar integrands is based on the comb component of the multi-particle solutions and the $\ell$-loop kernel. These two objects will be defined later. Based on these two objects, together with generalized BG currents, we will obtain the $\ell$-loop planar integrands recursively. We will also consider the Yang-Mills case and show some examples. It is worth mentioning that the $\ell$-loop planar integrands of the Yang-Mills theory can also be derived using planar variables [23, 24]. This method can be generalized to any theory to obtain the $\ell$-loop planar integrands by considering the planar component of the multi-particle solution to rule out the non-planar part. In this letter, the word “ $\ell$-loop integrands” is equivalent to “$\ell$-loop planar integrands” since we will not discuss the non-planar ones.
This letter is organized as follows. We will focus on the bi-adjoint scalar theory and the Yang-Mills theory. We first consider the bi-adjoint scalar theory and define the comb component of the multi-particle solutions and the $\ell$-loop kernels. Then we will show how to reach the $\ell$-loop integrands of bi-adjoint theory using our systematic approach and show some examples. Finally, we will show how to apply our method to the Yang-Mills theory. We will choose some Feynman diagrams as examples and show that our method can reproduce them.
@@ II Bi-adjoint: comb component and 1-loop kernel
where $\phi=\phi_{a\tilde{a}}T^{a}\otimes\tilde{T}^{\tilde{a}}$. The notation [[]]delimited-[][[]][ [ ] ] denotes the double commutator of two different generator $T^{a}$ and $\tilde{T}^{\tilde{a}}$, i.e.
@@
Using the perturbiner method [5], we can obtain the multi-particle solution of the equation of motion above:
@@
The next step is to define an $m$-way 1-loop kernel. We consider the comb component $\phi^{\rm comb}_{l12\cdots m|l12\cdots m}$, then using the sewing procedure [19] $\phi_{l|l}\to 1/l_{1}^{2}$ we can obtain the $m$-way 1-loop kernel
@@
$I^{\rm 1-loop}(12\cdots n|Q)=\sum_{m=2}^{n}\sum_{\text{ $m$-division of $P$}}I^{\rm kernel}_{1,m}\bigg{|}_{k_{i}\to k_{P_{i}}\atop\phi_{i|i}\to\Phi_{P_{i}|Q_{i}}}$ (8)
@@ III Graph factor
Let us see an example. In FIG. 1, each diagram has two loops. After ignoring the external legs and the corresponding vertices, there are only 2 propagators in the largest loop of each diagram. It is not hard to find that the values of both loops are 0 in each diagram. Hence, for these diagrams, we have $r=0$. Actually, only when the number of loops is greater than or equal to 3 will $r$ be possible to be nonzero.
In addition, there is an extra contribution to the symmetry factor. For 2-way kernels, it has a symmetry under flipping the diagram. By the word “flip”, we mean “turn upside down”. This will lead to some overcounts when we do the recursion introduced in the next section. Hence, we will regard the 2-way kernel, with all loops in it assigned a label, and its flipping as two different diagrams when recursion, and these two diagrams will both get an extra symmetry factor of 1/2, unless the position of labels of the 2-way kernel and its flipping coincide. We will see an example later.
@@ IV Bi-adjoint: Towards $\ell$-loop integrands
    Then, for $k$-divisions, which means we divide a set into $k$ parts, of each set, we set: i) $b_{n}$ itself to be one part of the $k$-division, ii) For a part not involving $a_{n}$, there can only be 1 element in the part, like $(b_{n}|a_{n}1|2|3)$. There is only 1 case for a given $k$ and a given set.
@@
$I^{\rm\ell-loop}(12\cdots n|Q)=\sum_{k=2}^{n}\sum_{\text{ $k$-division of $P$}}I^{\rm kernel}_{\ell,k}\bigg{|}_{k_{i}\to k_{P_{i}}\atop\phi_{i|i}\to\Phi_{P_{i}|Q_{i}}}$ (18)
@@
with $P=123\cdots n$. Note that the factor $m/\ell$ in the second line is very important to avoid overcounting. For example, consider a 3-loop integrand with 3 separate 1-loop parts. In our formula, it will give a factor of 3 since all the 1-loops are indistinguishable and we must add a factor of 1/3 to cancel this overcounting in the 1-loop kernel term. However, there are also some diagrams with a 1-loop part and a 2-loop part; this kind of diagram, based on the 1-loop kernel, is not overcounted but also receives a factor of 1/3. Then we must add a factor of 2/3 to the 2-loop kernel terms to cancel the factor of 1/3. Generalizing this to the $\ell$-loop case, we will get a factor of $m/\ell$ in front of the $m$-loop kernel terms.
The symmetry factors will not bother us in (LABEL:n-loop). These factors only affect the recursion of the loop kernel. Adding $\ell$-loop current $\Phi^{(\ell)}_{P|Q}$ to loop kernels does nothing to the symmetry factor.
@@
$I^{\rm 2-loop}(12|12)=\sum_{\text{ $2$-division of $(12)$}}I^{\rm kernel}_{2,2}\bigg{|}_{k_{i}\to k_{P_{i}}\atop\phi_{i|i}\to\Phi_{P_{i}|Q_{i}}}$ (19)
+12($+\frac{1}{2}(\sum_{\ell_{1},\ell_{2}\atop\ell_{1}+\ell_{2}=1}\sum_{\text{ $2$-division of $(12)$}}I^{\rm kernel}_{1,2}\bigg{|}_{k_{i}\to k_{P_{i}}\atop\phi_{i|i}\to\Phi^{(\ell_{i})}_{P_{i}|Q_{i}}})$
@@
$I^{\rm 2-loop}(12|12)=I^{\rm kernel}_{2,2}\bigg{|}_{\phi_{i|i}\to 1}$ (20)
@@
Note that a 2-way 1-loop kernel has a symmetry factor of 1/2, which is the origin of $1/4$ in the last term. We also give some more examples in the Supplementary Material.
V Yang-Mills: Some generalizations and examples
@@
We will use the Feynman gauge for the off-shell multi-particle solution $A_{P}$. The classical equation of motion of the Yang-Mills theory is
@@
Based on this, we have the 2-way 1-loop kernel by the replacement $A_{l\mu}A_{3\nu}\to\eta_{\mu\nu}/l_{1}^{2}$ in $A_{l12}\cdot A_{3}$ [19]:
@@
where $k_{l}=l_{1}$ and the factor of 1/2 is the symmetry factor. Note that here the last two terms correspond to the loop with a 4-point vertex. For the higher-loop kernels, one can construct them using the method in section III. There is another subtlety due to the appearance of 4-point vertices. For higher-way (more than 2) 1-loop kernels, we need to add an extra term besides the comb component in order to keep the cyclic invariance of the loop kernels. This term, called the contact component, is a component of the comb component, in which we assign the last deconcatenation as a 3-deconcatenation. For example, the contact component of $A_{1234\mu}$ is
@@
In the case $P=1234$, we have ignored all the 3-deconcatenation terms like $P=(12,3,4)$ since the last deconcatenation is a 2-deconcatenation like 12=(1,2)121212=(1,2)12 = ( 1 , 2 ). To construct $m$-way 1-loop kernels with $m>2$, we need to consider
$A^{\rm comb}_{l12\cdots m\mu}+A^{\rm contact}_{ln12\cdots(n-1)\mu}$ (27)
instead of only the comb component. After figuring out the graph factors similar to what we have done for the bi-adjoint scalars, a systematic approach to $\ell$-loop Yang-Mills integrands is established.
@@
Figure 3: 2-point 2-loop diagram 2
Now we will give some examples of 2-loop Feynman diagrams. Here, to simplify the calculation, we will just consider one of the possible divisions that can reach the diagrams and ignore some extra graph factors. We just aim to show that our method is consistent with the results from the Feynman rules. Consider the 2-point 2-loop diagram 1 (FIG. 2), we can generate this diagram by considering the last two terms of $I_{1,2}^{\text{kernel}}$, and then consider the 2-division $(b_{2}|a_{2},1,2)$ and the comb component $A^{\rm comb}_{a_{2}12}$. This will lead to the following result
$I^{\rm 2-loop}_{\rm FIG.1}(12)=$ (28)
$(1-d)\frac{(2l_{2}^{2}-2k_{2}\cdot l_{2})+(4d-6)(l_{2}\cdot e_{1})(l_{2}\cdot e_{2})}{2l_{1}^{2}l_{2}^{4}(l_{2}+k_{1})^{2}}.$
Note that the on-shell condition is $A_{i\mu}=\epsilon_{i\mu}$ with $\epsilon_{i}\cdot k_{i}=0$. A more non-trivial case is the 2-point 2-loop diagram 2 (FIG. 3), which corresponds to the terms with a propagator in the 1-loop kernel. We put the result in the Supplementary Material. Our results from these two diagrams match with the results from FeynCalc [28].
VI Conclusion and Outlook
In this letter, we give a systematic approach to $\ell$-loop planar integrands. We divide the $\ell$-loop planar integrands into 2 parts, the irreducible part and the reducible part. The irreducible part can be obtained from the $\ell$-loop kernel, which can be obtained recursively from the kernel with fewer loops, and some BG currents. The reducible part can be reached from the loop kernel with fewer loops and some generalized BG currents, which can be reached from loop integrands with fewer loops. Hence, the $\ell$-loop planar integrands can be determined by some objects with fewer loops, and all these constitute our recursion relation (LABEL:n-loop).
@@ Acknowledgements.
Supplementary Material
The 2-loop kernel with more ways
@@
    $(b_{2}|a_{2},1,2,3,4)+\text{cyclic}(1234)$ (38)
    $I^{(2)}_{2,4}=\frac{\phi_{1|1}\phi_{2|2}\phi_{3|3}\phi_{4|4}}{2l_{1}^{2}(l_{1}-l_{2})^{2}l_{2}^{4}(l_{2}+k_{1})^{2}(l_{2}+k_{1}+k_{2})^{2}(l_{2}+k_{1}+k_{2}+k_{3})^{2}}+\text{cyclic}(k_{1},k_{2},k_{3},k_{4})$ (39)
@@
    $I^{(3)}_{2,4}=\frac{\phi_{1|1}\phi_{2|2}\phi_{3|3}\phi_{4|4}}{2l_{1}^{2}(l_{1}-l_{2})^{2}(l_{1}+k_{1}+k_{2}+k_{2})^{2}l_{2}^{2}(l_{2}+k_{1})^{2}(l_{2}+k_{1}+k_{2})^{2}(l_{2}+k_{1}+k_{2}+k_{3})^{2}}+\text{cyclic}(k_{1},k_{2},k_{3},k_{4})$ (41)
@@
    $I^{(4)}_{2,4}=\frac{\phi_{1|1}\phi_{2|2}\phi_{3|3}\phi_{4|4}}{2l_{1}^{2}(l_{1}-l_{2})^{2}(l_{1}+k_{1}+k_{2})^{2}(l_{1}+k_{1}+k_{2}+k_{3})^{2}l_{2}^{2}(l_{2}+k_{1})^{2}(l_{2}+k_{1}+k_{2})^{2}}+\text{cyclic}(k_{1},k_{2},k_{3},k_{4})$ (43)
@@
    $I^{(5)}_{2,4}=\frac{\phi_{1|1}\phi_{2|2}\phi_{3|3}\phi_{4|4}}{2l_{1}^{2}(l_{1}-l_{2})^{2}(l_{1}+k_{1})^{2}(l_{1}+k_{1}+k_{2})^{2}(l_{1}+k_{1}+k_{2}+k_{3})^{2}l_{2}^{2}(l_{2}+k_{1})^{2}}+\text{cyclic}(k_{1},k_{2},k_{3},k_{4})$ (45)
@@
    $I^{(6)}_{2,4}=\frac{\phi_{1|1}\phi_{2|2}\phi_{3|3}\phi_{4|4}}{2l_{1}^{4}(l_{1}-l_{2})^{2}(l_{1}+k_{1})^{2}(l_{1}+k_{1}+k_{2})^{2}(l_{1}+k_{1}+k_{2}+k_{3})^{2}l_{2}^{2}}+\text{cyclic}(k_{1},k_{2},k_{3},k_{4})$ (47)
@@
$I^{\rm 2-loop}(123|123)=I^{\rm kernel}_{2,2}\bigg{|}_{k_{2}\to k_{23}\atop\phi_{1,1}\to 1,\ \phi_{2|2}\to\Phi_{23|23}}+I^{\rm kernel}_{2,2}\bigg{|}_{k_{1}\to k_{31}\atop\phi_{2,2}\to 1,\ \phi_{1|1}\to\Phi_{31|31}}+I^{\rm kernel}_{2,2}\bigg{|}_{k_{1}\to k_{12},\ k_{2}\to k_{3}\atop\phi_{2,2}\to 1,\ \phi_{1|1}\to\Phi_{12|12}}+I^{\rm kernel}_{2,3}\bigg{|}_{\phi_{i|i}\to 1}$ (49)
@@
The 2-loop gluon diagram example
For the 2-point 2-loop diagram 2 (FIG. 3), we consider the 2-division $(b_{2}|a_{2},1,2)$ and the comb component $A^{\rm comb}_{a_{2}12}$ to construct the corresponding off-shell diagram, and take the on-shell limit. The final result is
$2l_{1}^{2}(l_{1}+l_{2}+k_{1})^{2}l_{2}^{2}(l_{2}+k_{1})^{2}(l_{2}-k_{2})^{2}I^{\rm 2-loop}_{\rm FIG.2}(12)=$ (50)
$e_{1}\cdot e_{2}\bigg{(}(4d-6)\left(k_{2}\cdot l_{1}\right){}^{2}+2k_{2}\cdot l_{1}\left((4d-6)l_{1}\cdot l_{2}+(2d-5)l_{2}\cdot l_{2}+2k_{2}\cdot l_{2}\right)-2l_{2}\cdot l_{2}\left((1-2d)l_{1}\cdot l_{2}+15k_{2}\cdot l_{2}-2l_{1}\cdot l_{1}\right)$
$+2(2d-3)\left(l_{1}\cdot l_{2}\right){}^{2}+(d+4)\left(l_{2}\cdot l_{2}\right){}^{2}+20\left(k_{2}\cdot l_{2}\right){}^{2}-4\left(l_{1}\cdot l_{1}+l_{1}\cdot l_{2}\right)k_{2}\cdot l_{2}\bigg{)}+2(2d-3)e_{1}\cdot l_{1}\bigg{(}e_{2}\cdot l_{1}\left(l_{2}\cdot l_{2}-4k_{2}\cdot l_{2}\right)$
$+e_{2}\cdot l_{2}\left(3k_{2}\cdot l_{1}+k_{2}\cdot l_{2}-3l_{1}\cdot l_{2}-l_{2}\cdot l_{2}\right)\bigg{)}+e_{1}\cdot l_{2}\bigg{(}e_{2}\cdot l_{2}\left((19d-24)l_{2}\cdot l_{2}-2(2d-3)\left(3k_{2}\cdot l_{1}+10k_{2}\cdot l_{2}-6l_{1}\cdot l_{1}-3l_{1}\cdot l_{2}\right)\right)$
$+2(2d-3)e_{2}\cdot l_{1}\left(3k_{2}\cdot l_{1}+k_{2}\cdot l_{2}-3l_{1}\cdot l_{2}-l_{2}\cdot l_{2}\right)\bigg{)}$
Note that here we have done a transformation $l_{2}\to-l_{2}-k_{1}$ so that the result is easy to compare with the result from FeynCalc.

### v4

@@ Abstract
In this paper, we present a recursive method for $\ell$-loop planar integrands in colored quantum field theories. We start with the classical equation of motion and then pick out the comb component, which will help us to define the loop kernels. Then we construct the $\ell$-loop integrands based on some recursion rules for the $\ell$-loop kernels. Finally, we reach a recursion formula for the $\ell$-loop planar integrands. Our method can be easily generalized to general quantum field theories, even non-Lagrangian theories, to obtain the planar part of the whole $\ell$-loop integrands.
@@ I Introduction
Most of the development on the analytic side comes from the success of the on-shell methods, see [1, 2] for some reviews. Besides the on-shell method, off-shell methods have also been developed in various fields [3]. Off-shell objects have a richer structure and can help us know more about different theories. In addition, off-shell objects are very useful for generating loop integrands or computing amplitudes recursively. One of the most successful off-shell objects is the Berends-Giele (BG) current [4]. The BG current, which has only one off-shell leg, has many applications in different cases [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Hence, to develop off-shell techniques is very important both in theory and practical calculations.
@@
In this paper, we want to develop techniques based on the multi-particle solutions. We start with the equation of motion of the bi-adjoint theory, and finally reach the whole $\ell$-loop planar integrand results, which correspond to the partial amplitudes associated with a single-trace structure in the color-trace decomposition [1]. This systematic approach to $\ell$-loop planar integrands is based on the comb component of the multi-particle solutions and the $\ell$-loop kernel. These two objects will be defined later. Based on these two objects, together with generalized BG currents, we will obtain the $\ell$-loop planar integrands recursively. We will also consider the Yang-Mills case and show some examples. It is worth mentioning that the $\ell$-loop planar integrands of the Yang-Mills theory can also be derived using planar variables [24, 25]. This method can be generalized to any theory to obtain the $\ell$-loop planar integrands by considering the planar component of the multi-particle solution to rule out the non-planar part. In this paper, the word “ $\ell$-loop integrands” is equivalent to $\ell$-loop planar integrands” since we will not discuss the non-planar ones.
This paper is organized as follows. We will focus on the bi-adjoint scalar theory and the Yang-Mills theory. We first consider the bi-adjoint scalar theory and define the comb component of the multi-particle solutions and the $\ell$-loop kernels. Then we will show how to reach the $\ell$-loop integrands of bi-adjoint theory using our systematic approach and show some examples. Finally, we will show how to apply our method to the Yang-Mills theory. We will choose some Feynman diagrams as examples and show that our method can reproduce them.
@@ II Bi-adjoint: comb component and 1-loop kernel
where $\phi=\phi_{a\tilde{a}}T^{a}\otimes\tilde{T}^{\tilde{a}}$. The notation $[[]]$ denotes the double commutator of two different generator $T^{a}$ and $\tilde{T}^{\tilde{a}}$, i.e.
@@
Using the perturbiner method [26, 27, 28, 29, 5], we can obtain the multi-particle solution of the equation of motion above:
@@
The next step is to define an $m$-way 1-loop kernel. We consider the comb component with the last propagator $s_{12\cdots m}\phi^{\rm comb}_{l12\cdots m|l12\cdots m}$, then using the sewing procedure [21] $\phi_{l|l}\to 1/l_{1}^{2}$ and imposing the momentum conservation for loop integrands, we can obtain the $m$-way 1-loop kernel
@@
[…]
@@ III Graph factor
In addition, there is an extra contribution to the symmetry factor. For 2-way kernels, there is a symmetry under flipping the diagram. By the word “flip”, we mean “turn upside down”. This will lead to some overcounts when we do the recursion introduced in the next section. Hence, we will regard the 2-way kernel and its flipping as two different diagrams when recursion, and these two diagrams will both get an extra symmetry factor of 1/2, unless the 2-way kernel and its flipping coincide. We will see an example later. Actually, in the bi-adjoint scalar case, all symmetry factors come from the 2-way kernel parts of the Feynman diagrams. Every 2-way kernel part contributes a factor of 1/2. A more elegant way to deal with it in the bi-adjoint theory has been demonstrated in [30].
@@ IV Bi-adjoint: Towards $\ell$-loop integrands
    Then, for $k$-divisions, which means we divide a set into $k$ parts, of each set, we set: i) $b_{\ell}$ itself to be one part of the $k$-division, ii) For a part not involving $a_{\ell}$, there can only be 1 element in the part, like $(b_{\ell}|a_{\ell},1|2|3)$. There is only 1 case for a given $k$ and a given set.
@@
[…]
@@
with $P=123\cdots n$. Note that the factor $m/\ell$ in the second line is very important to avoid overcounting. Consider a given reducible diagram, it has $n_{m}$ number of $m$-loop parts and it satisfies $\sum_{m=1}^{\ell-1}m\times n_{m}=\ell$. For the $m$-loop kernel term in the second line of (LABEL:n-loop), there will be $n_{m}$ number of overcounts since every $m$-loop part is equivalent. After adding the factor $m/\ell$, the total counting is $\sum_{m=1}^{\ell-1}n_{m}\times m/\ell=1$. Hence we have canceled all the overcounts. The symmetry factors will not bother us in (LABEL:n-loop). These factors only affect the recursion of the loop kernel. Adding $\ell$-loop current $\Phi^{(\ell)}_{P|Q}$ to loop kernels does nothing to the symmetry factor.
@@
[…]
$+\frac{1}{2}(\sum_{\ell_{1},\ell_{2}\atop\ell_{1}+\ell_{2}=1}\sum_{\text{ $2$-division of $(12)$}}I^{\rm kernel}_{1,2}\bigg|_{k_{i}\to k_{P_{i}}\atop\phi_{i|i}\to\Phi^{(\ell_{i})}_{P_{i}|Q_{i}}})$
@@
$1\displaystyle I^{\rm 2-loop}(2|2)=I^{\rm kernel}_{2,2}\bigg|_{\phi_{i|i}\to 1}$ (20)
@@
Note that a 2-way 1-loop kernel has a symmetry factor of 1/2, which is the origin of $1/4$ in the last term. We also give some more examples in the appendix.
V Yang-Mills: A more general case
@@
We will use the Feynman gauge for the off-shell multi-particle solution of gluons. The classical equation of motion of the Yang-Mills theory is
@@
Based on this, we obtain the 2-way 1-loop kernel after the replacement $A_{l\mu}A_{3\nu}\to\eta_{\mu\nu}/l_{1}^{2}$ in $A_{l12}\cdot A_{3}$ [21]:
@@
where $k_{l}=l_{1}$ and the factor of 1/2 is the symmetry factor.
Note that here the last two terms correspond to the loop with a 4-point vertex. For the higher-loop kernels, one can construct them using the method in section III. There is another subtlety due to the appearance of 4-point vertices. For higher-way (more than 2) 1-loop kernels, if we do the same thing as the bi-adjoint theory case, i.e., consider $A^{\rm comb}_{l12\cdots m\mu}$ only, we will find that there must be a propagator between leg 1 and leg $m$. However, there also exists a case that leg 1 and leg $m$ share a common 4-vertex. The lack of this case will break the cyclic invariance of the 1-loop kernels. Hence, we need to add an extra term besides the comb component in order to keep the cyclic invariance of 1-loop kernels. This term, called the contact component, has the same definition as the comb component but with the last deconcatenation being a 3-deconcatenation. For example, the contact component of $A_{1234\mu}$ is
@@
In the case $P=1234$, we have ignored all the 3-deconcatenation terms like $P=(12,3,4)$ since the last deconcatenation is a 2-deconcatenation like $12=(1,2)$. To construct $m$-way 1-loop kernels with $m>2$, we need to consider
$A^{\rm comb}_{l12\cdots m\mu}+A^{\rm contact}_{lm12\cdots(m-1)\mu}$ (27)
instead of only the comb component. Such contact components appear only when we construct the 1-loop kernels. When we go to higher-loop integrands recursively, we do not need to consider contact components anymore. After figuring out the graph factors similar to what we have done for the bi-adjoint scalars, a systematic approach to $\ell$ gluon-loop Yang-Mills integrands is established. Note that the contribution from ghost-loops can be restored by considering the equation of motion involving ghost fields [21].
VI Yang-Mills: Examples
In this section, we will give some examples of 2-loop YM planar integrands. Here, to simplify the calculation, we will just consider one of the possible divisions that can reach given Feynman diagrams and ignore some extra graph factors. We just aim to show that our method is consistent with the results from the Feynman rules. Consider the 2-point 2-loop diagram 1 (FIG. 2), we can generate this diagram by considering the last two terms of $I_{1,2}^{\text{kernel}}$ for the YM theory, and then consider the 2-division $(b_{2}|a_{2},1,2)$ and the comb component $A^{\rm comb}_{a_{2}12\mu}$. This will lead to the following result
[…]
Note that we have used the on-shell condition here.
For the 2-point 2-loop diagram 2 (FIG. 3), we consider the 2-division $(b_{2}|a_{2},1,2)$ and the comb component $A^{\rm comb}_{a_{2}12\mu}$ again, together with all but the last two terms of $I_{1,2}^{\rm kernel}$, to construct the corresponding off-shell diagram, and then take the on-shell limit. The final result is
[…]
$\epsilon_{1}\cdot\epsilon_{2}\bigg((4d-6)\left(k_{2}\cdot l_{1}\right){}^{2}+2k_{2}\cdot l_{1}\left((4d-6)l_{1}\cdot l_{2}+(2d-5)l_{2}\cdot l_{2}+2k_{2}\cdot l_{2}\right)-2l_{2}\cdot l_{2}\left((1-2d)l_{1}\cdot l_{2}+15k_{2}\cdot l_{2}-2l_{1}\cdot l_{1}\right)$
[…]
$+\epsilon_{2}\cdot l_{2}\left(3k_{2}\cdot l_{1}+k_{2}\cdot l_{2}-3l_{1}\cdot l_{2}-l_{2}\cdot l_{2}\right)\bigg)+\epsilon_{1}\cdot l_{2}\bigg(\epsilon_{2}\cdot l_{2}\left((19d-24)l_{2}\cdot l_{2}-2(2d-3)\left(3k_{2}\cdot l_{1}+10k_{2}\cdot l_{2}-6l_{1}\cdot l_{1}-3l_{1}\cdot l_{2}\right)\right)$
$+2(2d-3)\epsilon_{2}\cdot l_{1}\left(3k_{2}\cdot l_{1}+k_{2}\cdot l_{2}-3l_{1}\cdot l_{2}-l_{2}\cdot l_{2}\right)\bigg)$
Note that here we have done a shift $l_{2}\to-l_{2}-k_{1}$ so that the result is easy to compare with the result from FeynCalc. Both our results match the results from FeynCalc [34].
@@ VI Yang-Mills: Examples
Figure 3: 2-point 2-loop diagram 2 ($l_{2}$ here is the shifted one)
VII Discussions
In this paper, we give a systematic approach to $\ell$-loop planar integrands. We divide the $\ell$-loop planar integrands into 2 parts, the irreducible part and the reducible part. The irreducible part can be obtained from the $\ell$-loop kernel, which can be obtained recursively from the kernel with fewer loops, and some BG currents. The reducible part can be reached from the loop kernel with fewer loops and some generalized BG currents, which can be reached from loop integrands with fewer loops. Hence, the $\ell$-loop planar integrands can be determined by some objects with fewer loops, and all these constitute our recursion relation (LABEL:n-loop).
This new recursion method, which starts with the classical equation of motion, is somehow a higher-loop version of the Berends-Giele recursion, which organizes the Feynman rules in a proper way to make the recursion properties manifest. In our $\ell$-loop recursion, we also creatively organized the Feynman rules so that we can reach $\ell$-loop integrands recursively. In contrast to other methods towards $\ell$-loop planar integrands, like [25, 24], which can only be used in limited theories, our method has broader applications since it is based on the Feynman rules rather than some special properties of some specific theories.
@@ Acknowledgements.
Appendix A The 2-loop kernel with more ways
@@
[…]
    $I^{(2)}_{2,4}=\frac{\phi_{1|1}\phi_{2|2}\phi_{3|3}\phi_{4|4}}{l_{1}^{2}(l_{1}-l_{2})^{2}l_{2}^{4}(l_{2}+k_{1})^{2}(l_{2}+k_{1}+k_{2})^{2}(l_{2}+k_{1}+k_{2}+k_{3})^{2}}+\text{cyclic}(k_{1},k_{2},k_{3},k_{4})$ (40)
@@
    $I^{(3)}_{2,4}=\frac{\phi_{1|1}\phi_{2|2}\phi_{3|3}\phi_{4|4}}{l_{1}^{2}(l_{1}-l_{2})^{2}(l_{1}+k_{1}+k_{2}+k_{3})^{2}l_{2}^{2}(l_{2}+k_{1})^{2}(l_{2}+k_{1}+k_{2})^{2}(l_{2}+k_{1}+k_{2}+k_{3})^{2}}+\text{cyclic}(k_{1},k_{2},k_{3},k_{4})$ (42)
@@
    $I^{(4)}_{2,4}=\frac{\phi_{1|1}\phi_{2|2}\phi_{3|3}\phi_{4|4}}{l_{1}^{2}(l_{1}-l_{2})^{2}(l_{1}+k_{1}+k_{2})^{2}(l_{1}+k_{1}+k_{2}+k_{3})^{2}l_{2}^{2}(l_{2}+k_{1})^{2}(l_{2}+k_{1}+k_{2})^{2}}+\text{cyclic}(k_{1},k_{2},k_{3},k_{4})$ (44)
@@
    $I^{(5)}_{2,4}=\frac{\phi_{1|1}\phi_{2|2}\phi_{3|3}\phi_{4|4}}{l_{1}^{2}(l_{1}-l_{2})^{2}(l_{1}+k_{1})^{2}(l_{1}+k_{1}+k_{2})^{2}(l_{1}+k_{1}+k_{2}+k_{3})^{2}l_{2}^{2}(l_{2}+k_{1})^{2}}+\text{cyclic}(k_{1},k_{2},k_{3},k_{4})$ (46)
@@
    $I^{(6)}_{2,4}=\frac{\phi_{1|1}\phi_{2|2}\phi_{3|3}\phi_{4|4}}{l_{1}^{4}(l_{1}-l_{2})^{2}(l_{1}+k_{1})^{2}(l_{1}+k_{1}+k_{2})^{2}(l_{1}+k_{1}+k_{2}+k_{3})^{2}l_{2}^{2}}+\text{cyclic}(k_{1},k_{2},k_{3},k_{4})$ (48)
@@
$1\displaystyle I^{\rm 2-loop}(23|23)=I^{\rm kernel}_{2,2}\bigg|_{k_{2}\to k_{23}\atop\phi_{1,1}\to 1,\ \phi_{2|2}\to\Phi_{23|23}}+I^{\rm kernel}_{2,2}\bigg|_{k_{1}\to k_{31}\atop\phi_{2,2}\to 1,\ \phi_{1|1}\to\Phi_{31|31}}+I^{\rm kernel}_{2,2}\bigg|_{k_{1}\to k_{12},\ k_{2}\to k_{3}\atop\phi_{2,2}\to 1,\ \phi_{1|1}\to\Phi_{12|12}}+I^{\rm kernel}_{2,3}\bigg|_{\phi_{i|i}\to 1}$ (50)